OpenStudy (anonymous):

Write the standard form of the equation of the parabola that has the indicated vertex and whose graph passes through the given point. Vertex: (−1, −2); point: (0, 0)

OpenStudy (anonymous):

f(x)=a(0+1)^2-2

OpenStudy (anonymous):

y = (x+1)^2 - 2 y = x^2 + 2x + 1 - 2 y = x^2 + 2x -1